OpenStudy (anonymous):

A plane landing on a small tropical island has just 64 m of runway on which to stop. If its initial speed is 66 m/s, what is the maximum acceleration of the plane during landing, assuming it to be constant? Answer in units of m/s 2 How long does it take for the plane to stop with this acceleration? Answer in units of s

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@Agent47 help

OpenStudy (anonymous):

This looks familiar. Didn't you already do this?

OpenStudy (anonymous):

i got it wrong

OpenStudy (anonymous):

well the second part

OpenStudy (anonymous):

OK, I'll take a closer look at it and see if I can show you the setup in a clear way. Give me a few seconds.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The acceleration you got last time, -34.03125 m/s^2 is correct, and I'm getting 64/33 s for the time to stop using two different formulas.

OpenStudy (anonymous):

yea i got -34.03125 but whats the 64/33

OpenStudy (anonymous):

64/33 ≈ 1.94s

OpenStudy (anonymous):

Though that should be rounded to 1.9s given the limits of precision you have.

OpenStudy (anonymous):

that answer was wrong

OpenStudy (anonymous):

hmmm.. The acceleration is found by: \[(v_f)^2-(v_o)^2=2a(x_f-x_o)\] \[(0m/s)^2-(66m/s)^2=2a(64m-0m) \rightarrow\] \[a=\frac{(0m/s)^2-(66m/s)^2}{2(64m-0m)} \space \space = -34.03125m/s^2\]

OpenStudy (anonymous):

The time can be found by either \[x_f=x_o+v_ot+\frac{1}{2}at^2 \space \space or \space \space v_f-v_o=at\]

OpenStudy (anonymous):

the time part confuses me

OpenStudy (anonymous):

\[64m=0m+66m/s(t)+\frac{1}{2}(-34.03125m/s^2)(t)^2 \space \space or\] \[0m/s−66m/s=(-34.03125m/s^2)(t)\]

OpenStudy (anonymous):

Using either of those equations, \[t=1\frac{31}{33}≈1.939393...\]

OpenStudy (anonymous):

divide -66 by -34.01325 and that equals t. time

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

1.939393

OpenStudy (anonymous):

thanks i got the answer by myself the first time but i dont know what happend thank you very much

OpenStudy (anonymous):

It all checks out. I don't know why you're getting errors. Personally, with all the uncertainty, I would round that up to 2 seconds.

OpenStudy (anonymous):

yea thanks i know now